OpenStudy (rcp031):

Three people, P, Q and R contribute to a fund. P provides 3/5 of the total, Q provides 2/3 of the remainder, and R provides £8. Determine (a) the total of the fund, (b) the contributions of P and Q. [(a) £60 (b) £36, £16]

OpenStudy (anonymous):

If P gives 3/5 of the total the remaining is 2/5.( coz the whole thing is 5/5. subtract 3/5 frm 5/5. u get 2/5) Q gives 2/3 of that remainder. that is, 2/3 of 2/5. [of = multiplication. that means he gave 2/3 x 2/5] Amount Q gave = 4/15 Then if Q gave 2/3 of the remainder, that means the rest which is 1/3 has to be given by R. So 1/3 of 2/5. R gave = 1/3 x 2/5 = 2/15. It's given that 2/15 = £ 8. 2/15 ----------> £8 1/15 ----------> £8/2 = £ 4 amount Q gave = 4/15 --------> 4 x 4 = £16 amount P gave = 3/5 = 9/15 ------------> 4 x 9 = £36. total = £ (36 + 8 + 16) = £ 60